# Finite-Time Analysis of Simultaneous Double Q-learning

### v1

@@ 6.1 Overall plans
The overall flow of the proposed analysis is given in Figure 3. The texts highlighted with blue indicate the dynamic systems we will deal with for our analysis. The red arrows represent the directions we will follow for the proof. The overall process is summarized as follows: Step 1: The finite-time error bound of $Q_{k}^{\text{err}_{UL}}$ is obtained by using its corresponding linear system structure. Then, based on the error bound on $Q_{k}^{\text{err}_{\text{UL}}}$, the finite-time error bound on $Q_{k}^{\text{err}_{U}}$ can be derived. Step 2: following similar lines as in Step 1, one can derive the error bound on $Q_{k}^{\text{err}}$ based on the error bound on $Q_{k}^{\text{err}_{U}}$ and $Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{L}}$. Step 3: Using the error bound on $Q_{k}^{\text{err}}$ and the linear structures of $Q_{k}^{A_{L}}-Q^{*}$ and $Q_{k}^{B_{L}}-Q^{*}$, the finite-time error bounds on $Q_{k}^{A_{L}}-Q^{*}$ and $Q_{k}^{B_{L}}-Q^{*}$ can be derived. Step 4: By obtaining a subtraction system which can be obtained by subtracting the error lower comparison system from the error upper comparison system, the convergence of $Q_{k}^{A_{U}}-Q^{*}$ and $Q_{k}^{B_{U}}-Q^{*}$ can be shown. Step 5: Using the previous results, we can obtain a finite-time error bound on the iterates of SDQ. These steps will be detailed in Appendix.
Figure 3: Overall flow of the proposed analysis
7 Discussion and Conclusion
In this paper, we present a novel variant of double $Q$-learning, called simultaneous double $Q$-learning (SDQ), which empirically achieves faster convergence than the original double $Q$-learning by eliminating the random selection of double $Q$-estimators. Moreover, SDQ introduces a novel switching system perspective by swapping the roles of the two $Q$-estimators. By leveraging switching system representations, we establish novel finite-time expected error bounds for SDQ. It is also important to note that while the random $Q$-estimator selection step is eliminated in SDQ, it needs for the additional initialization step for the $Q$-estimator, which can affect the overall convergence properties. Therefore, future researches can aim to address this initialization effect.
Supplementary material
@@ Appendix A Convergence of stochastic linear system
To aid in understanding the intricacies of the convergence of simultaneous double $Q$-learning(SDQ), as explained in Sections B and C, let us consider the following stochastic linear system, which offers a helpful conceptual framework:
@@
where $x_{k}\in\mathbb{R}^{n}$ is the state, and $A$ is system matrix, and $\alpha v_{k}$ is the stochastic noise with a constant $\alpha\in(0,1)$. Here, we will first investigate a finite-time error analysis of the state of (15), and it will be used in the proof of SDQ. To this end, let us assume that the covariance matrix of $v_{k}$ is bounded as $\mathbb{E}[v_{k}v_{k}^{T}]\leq V_{\text{max}}$ and $V_{\text{max}}>0$. Then, it can be proved that the maximum eigenvalue of $\mathbb{E}[v_{k}v_{k}^{T}]$ can be bounded by $V_{\text{max}}$.
Lemma 1.
[13] The maximum eigenvalue of $\mathbb{E}[v_{k}v_{k}^{T}]$ is bounded as
$\lambda_{\text{max}}(\mathbb{E}[v_{k}v_{k}^{T}])\leq V_{\text{max}}$
for all $k\geq 0$, where $V_{\text{max}}>0$ is from our assumption.
@@
The system matrix $A$ satisfies $A\leq\rho$ with some constant $\rho\in(0,1)$.
@@
Lemma 2.
[13] We have the following bound:
$\text{tr}(X_{k})\leq\frac{9n^{2}\alpha}{d_{\text{min}}(1-\gamma)^{3}}+\lVert x_{0}\rVert^{2}_{2}n^{2}\rho^{2k}$
@@
$\begin{split}\lambda_{\text{max}}(X_{k})&\leq\alpha^{2}\sum_{i=0}^{k-1}\lambda_{\text{max}}(A^{i}V_{k-i-1}(A^{T})^{i})+\lambda_{\text{max}}(A^{k}X_{0}(A^{T})^{k})\\ &\leq\alpha^{2}\sup_{j\geq 0}\lambda_{\text{max}}\sum_{i=0}^{k-1}\lambda_{\text{max}}(A^{i}(A^{T})^{i})+\lambda_{\text{max}}(X_{0})\lambda_{\text{max}}(A^{k}(A^{T})^{k})\\ &=\alpha^{2}\sup_{j\geq 0}\lambda_{\text{max}}(V_{j})\sum_{i=0}^{k-1}\lVert A^{i}\rVert_{2}^{2}+\lambda_{\text{max}}(X_{0})\lVert A^{k}\rVert_{2}^{2}\\ &\leq\alpha^{2}V_{\text{max}}n\sum_{i=0}^{k-1}\lVert A^{i}\rVert_{\infty}^{2}+n\lambda_{\text{max}}(X_{0})\lVert A^{k}\rVert_{\infty}^{2}\\ &\leq\alpha^{2}V_{\text{max}}n\sum_{i=0}^{k-1}\rho^{2i}+n\lambda_{\text{max}}(X_{0})\rho^{2k}\\ &\leq\alpha^{2}V_{\text{max}}n\lim_{k\rightarrow\infty}\sum_{i=0}^{k-1}\rho^{2i}+n\lambda_{\text{max}}(X_{0})\rho^{2k}\\ &\leq\frac{\alpha^{2}V_{\text{max}}n}{1-\rho^{2}}+n\lambda_{\text{max}}(X_{0})\rho^{2k}\\ &\leq\frac{\alpha^{2}V_{\text{max}}n}{1-\rho}+n\lambda_{\text{max}}(X_{0})\rho^{2k}\end{split}$
where the first inequality is due to $A^{i}V_{k-i-1}(A^{T})^{i}\succeq 0$ and $A^{k}X_{0}(A^{T})^{k}\succeq 0$, the third inequality comes from Lemma 1, $\lVert\cdot\rVert_{2}\leq\sqrt{n}\lVert\cdot\rVert_{\infty}$, the fourth inequality is due to A.1, and the sixth and last inequalities come from $\rho\in(0,1)$. On the other hand, since $X_{k}\succeq 0$, the diagonal elements are nonnegative. Therefore, we have $\text{tr}(X_{k})\leq n\lambda_{\text{max}}(X_{k})$. Combining the last two inequalities leads to
@@
Theorem A.2.
[13] For any $k\geq 0$, we have
$\mathbb{E}\left[\lVert x_{k}\rVert_{2}\right]$ $\leq\frac{3\alpha^{1/2}n}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+n\lVert x_{0}\rVert_{2}\rho^{k}$ (16)
@@
$\begin{split}\mathbb{E}[\lVert x_{k}\rVert_{2}^{2}]=\mathbb{E}[\lVert x_{k}^{T}x_{k}\rVert]=\mathbb{E}[\lVert\text{tr}(x_{k}^{T}x_{k})\rVert]=\mathbb{E}[\lVert\text{tr}(x_{k}x_{k}^{T})\rVert]=\mathbb{E}[\text{tr}(X_{k})],\end{split}$
@@
$\mathbb{E}\biggl{[}\biggl{\lVert}A^{k}x_{0}+\sum_{j=0}^{k-1}\alpha A^{(k-1)-j}v_{j}\biggr{\rVert}_{2}\biggl{]}\leq\frac{3\alpha^{1/2}n}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+n\lVert x_{0}\rVert_{2}A^{k}$
@@ Appendix B Detailed analysis result of the convergence of SDQ ($Q_{k}^{\text{err}}$ part)
In this section, we present a detailed analysis process. To establish the groundwork for our proof, we first introduce an auxiliary Lemma demonstrating the boundedness of the system matrix.
Lemma 1.
[13] For any $Q\in{\mathbb{R}}^{\lvert S\rvert\lvert A\rvert}$, we have
@@
$\begin{split}\lVert A_{Q}\rVert_{\infty}&=\max_{i\in\{1,2,...,\lvert S\rvert\lvert A\rvert\}}{1+\alpha[D]_{ii}(\gamma-1)}\\ &=1-\alpha\min_{(s,a)\in\mathit{S}\times\mathit{A}}d(s,a)(1-\gamma)\\ &=\rho,\end{split}$
@@ B.1 Convergence of $Q_{k}^{\text{err}_{U}}$
$Q_{k+1}^{\text{err}_{U}}=(I+\alpha\gamma DP\Pi_{Q_{k}^{\text{err}_{U}}}-\alpha D)Q_{k}^{\text{err}_{U}}+\alpha w_{k}^{A}-\alpha w_{k}^{B},\quad Q_{0}^{\text{err}_{U}}\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert},$ (18)
@@
Proposition B.1.
Suppose $Q_{0}^{\text{err}_{U}}\geq Q_{0}^{\text{err}}$, where $\geq$” is used as the element-wise inequality. Then, we have
$Q_{k}^{\text{err}_{U}}\geq Q_{k}^{\text{err}}$
@@
$\begin{split}Q_{k+1}^{\text{err}}&=Q_{k}^{\text{err}}+\alpha\gamma DP\Pi_{Q_{k}^{B}}Q_{k}^{A}-\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{B}-\alpha DQ_{k}^{\text{err}}\\ &\quad+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &\leq Q_{k}^{\text{err}}+\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{A}-\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{B}-\alpha DQ_{k}^{\text{err}}\\ &\quad+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &=Q_{k}^{\text{err}}+\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{\text{err}}-\alpha DQ_{k}^{\text{err}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &=(I+\alpha\gamma DP\Pi_{Q_{k}^{A}}-\alpha D)Q_{k}^{\text{err}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &\leq(I+\alpha\gamma DP\Pi_{Q_{k}^{\text{err}}}-\alpha D)Q_{k}^{\text{err}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &\leq(I+\alpha\gamma DP\Pi_{Q_{k}^{\text{err}}}-\alpha D)Q_{k}^{\text{err}_{U}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &\leq(I+\alpha\gamma DP\Pi_{Q_{k}^{\text{err}_{U}}}-\alpha D)Q_{k}^{\text{err}_{U}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &=Q_{k+1}^{\text{err}_{U}},\end{split}$
@@
To prove the convergence of $Q_{k}^{\text{err}_{U}}$, we consider another comparison system $Q_{k}^{\text{err}_{UL}}$ which is a lower comparison system of $Q_{k}^{\text{err}_{U}}$. In the following subsection, a convergence analysis of $Q_{k}^{\text{err}_{UL}}$ is presented.
B.2 Convergence of $Q_{k}^{\text{err}_{UL}}$
Let us write the $Q_{k}^{\text{err}_{UL}}$, which has the following form:
$Q_{k+1}^{\text{err}_{UL}}=(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)Q_{k}^{\text{err}_{UL}}+\alpha w_{k}^{A}-\alpha w_{k}^{B},$ (19)
where the stochastic noises $w_{k}^{A}$ and $w_{k}^{B}$ are identical to those in the original system (5.1). Note that the system is the lower comparison system of the upper comparison system corresponding to $Q_{k}^{\text{err}}$.
In the following proposition, we prove that $Q_{k}^{\text{err}_{UL}}$ is a lower comparison system of $Q_{k}^{\text{err}_{U}}$.
Proposition B.2.

### v2

@@ 6.1 Overall plans
The overall flow of the proposed analysis is given in Figure 4. The texts highlighted with blue indicate the dynamic systems we will deal with for our analysis. The red arrows represent the directions we will follow for the proof. The overall process is summarized as follows: Step 1: The finite-time error bound of $Q_{k}^{\text{err}_{UL}}$ is obtained by using its corresponding linear system structure. Then, based on the error bound on $Q_{k}^{\text{err}_{UL}}$, the finite-time error bound on $Q_{k}^{\text{err}_{U}}$ can be derived. Step 2: following similar lines as in Step 1, one can derive the error bound on $Q_{k}^{\text{err}}$ based on the error bound on $Q_{k}^{\text{err}_{U}}$ and $Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{L}}$. Step 3: Using the error bound on $Q_{k}^{\text{err}}$ and the linear structures of $Q_{k}^{A_{L}}-Q^{*}$ and $Q_{k}^{B_{L}}-Q^{*}$, the finite-time error bounds on $Q_{k}^{A_{L}}-Q^{*}$ and $Q_{k}^{B_{L}}-Q^{*}$ can be derived. Step 4: By obtaining a subtraction system which can be obtained by subtracting the error lower comparison system from the error upper comparison system, the convergence of $Q_{k}^{A_{U}}-Q^{*}$ and $Q_{k}^{B_{U}}-Q^{*}$ can be shown. Step 5: Using the previous results, we can obtain a finite-time error bound on the iterates of SDQ. These steps will be detailed in Appendix.
7 Conclusion
In this paper, we present a novel variant of double $Q$-learning, called SDQ, which mitigates the maximization bias of standard $Q$-learning by using two separate $Q$-estimators and eliminating the random selection step. By alternating the roles of the two estimators, SDQ offers a novel switching system interpretation. Empirical results indicate that SDQ converges faster than the original double $Q$-learning. Based on this representation, we derive new finite-time expected error bounds that complement existing results. Future work will focus on tightening the dimensional dependence of the theoretical bound by developing refined analytical techniques that account for the coupled structure of the two estimators. We also plan to extend SDQ to function approximation and adaptive settings to further enhance convergence and robustness.
Acknowledgement
The work was supported by the Institute of Information Communications Technology Planning Evaluation (IITP) funded by the Korea government under Grant 2022-0-00469.
@@ Appendix A Convergence of stochastic linear system
To aid in understanding the intricacies of the convergence of SDQ, as explained in B and C, let us consider the following stochastic linear system, which offers a helpful conceptual framework:
@@
where $x_{k}\in\mathbb{R}^{n}$ is the state, and $A$ is system matrix, and $\alpha v_{k}$ is the stochastic noise with a constant $\alpha\in(0,1)$. Here, we will first investigate a finite-time error analysis of the state of (16), and it will be used in the proof of SDQ. To this end, let us assume that the noise energy of $v_{k}$ is bounded as $\mathbb{E}[v_{k}^{T}v_{k}]\leq V_{\text{max}}$ and $V_{\text{max}}>0$. Then, it can be proved that the maximum eigenvalue of $\mathbb{E}[v_{k}v_{k}^{T}]$ can be bounded by $V_{\text{max}}$.
Lemma 2 ([13]).
The maximum eigenvalue of $\mathbb{E}[v_{k}v_{k}^{T}]$ is bounded as
$\lambda_{\mathrm{max}}(\mathbb{E}[v_{k}v_{k}^{T}])\leq V_{\mathrm{max}}$
for all $k\geq 0$, where $V_{\mathrm{max}}>0$ is from our assumption.
@@
The system matrix $A$ satisfies $\|A\|_{\infty}\leq\rho$ for some constant $\rho\in(0,1)$.
@@
Lemma 3 ([13]).
We have the following bound:
$\mathrm{tr}(X_{k})\leq\frac{9n^{2}\alpha}{d_{\mathrm{min}}(1-\gamma)^{3}}+\lVert x_{0}\rVert^{2}_{2}n^{2}\rho^{2k}$
@@
$\begin{split}\lambda_{\text{max}}(X_{k})&\leq\alpha^{2}\sum_{i=0}^{k-1}\lambda_{\text{max}}(A^{i}V_{k-i-1}(A^{T})^{i})+\lambda_{\text{max}}(A^{k}X_{0}(A^{T})^{k})\\ &\leq\alpha^{2}\sup_{j\geq 0}\lambda_{\text{max}}(V_{j})\sum_{i=0}^{k-1}\lambda_{\text{max}}(A^{i}(A^{T})^{i})+\lambda_{\text{max}}(X_{0})\lambda_{\text{max}}(A^{k}(A^{T})^{k})\\ &=\alpha^{2}\sup_{j\geq 0}\lambda_{\text{max}}(V_{j})\sum_{i=0}^{k-1}\lVert A^{i}\rVert_{2}^{2}+\lambda_{\text{max}}(X_{0})\lVert A^{k}\rVert_{2}^{2}\\ &\leq\alpha^{2}V_{\text{max}}n\sum_{i=0}^{k-1}\lVert A^{i}\rVert_{\infty}^{2}+n\lambda_{\text{max}}(X_{0})\lVert A^{k}\rVert_{\infty}^{2}\\ &\leq\alpha^{2}V_{\text{max}}n\sum_{i=0}^{k-1}\rho^{2i}+n\lambda_{\text{max}}(X_{0})\rho^{2k}\\ &\leq\alpha^{2}V_{\text{max}}n\lim_{k\rightarrow\infty}\sum_{i=0}^{k-1}\rho^{2i}+n\lambda_{\text{max}}(X_{0})\rho^{2k}\\ &\leq\frac{\alpha^{2}V_{\text{max}}n}{1-\rho^{2}}+n\lambda_{\text{max}}(X_{0})\rho^{2k}\\ &\leq\frac{\alpha^{2}V_{\text{max}}n}{1-\rho}+n\lambda_{\text{max}}(X_{0})\rho^{2k}\end{split}$
where the first inequality is due to $A^{i}V_{k-i-1}(A^{T})^{i}\succeq 0$ and $A^{k}X_{0}(A^{T})^{k}\succeq 0$, the third inequality comes from Lemma 2, $\lVert\cdot\rVert_{2}\leq\sqrt{n}\lVert\cdot\rVert_{\infty}$, the fourth inequality is due to Assumption A.5, and the sixth and last inequalities come from $\rho\in(0,1)$. On the other hand, since $X_{k}\succeq 0$, the diagonal elements are nonnegative. Therefore, we have $\text{tr}(X_{k})\leq n\lambda_{\text{max}}(X_{k})$. Combining the last two inequalities leads to
@@
Theorem A.1 ([13]).
For any $k\geq 0$, we have
$\mathbb{E}\left[\lVert x_{k}\rVert_{2}\right]$ $\leq\frac{3\alpha^{1/2}n}{d_{\mathrm{min}}^{1/2}(1-\gamma)^{3/2}}+n\lVert x_{0}\rVert_{2}\rho^{k}.$ (17)
@@
$\begin{split}\mathbb{E}[\lVert x_{k}\rVert_{2}^{2}]=\mathbb{E}[x_{k}^{T}x_{k}]=\mathbb{E}[\text{tr}(x_{k}^{T}x_{k})]=\mathbb{E}[\text{tr}(x_{k}x_{k}^{T})]=\mathbb{E}[\text{tr}(X_{k})],\end{split}$
@@
$\mathbb{E}\biggl[\biggl\lVert A^{k}x_{0}+\sum_{j=0}^{k-1}\alpha A^{(k-1)-j}v_{j}\biggr\rVert_{2}\biggl]\leq\frac{3\alpha^{1/2}n}{d_{\mathrm{min}}^{1/2}(1-\gamma)^{3/2}}+n\lVert x_{0}\rVert_{2}\rho^{k}.$
@@ Appendix B Detailed analysis result of the convergence of SDQ ($Q_{k}^{\text{err}}$ part)
In this section, we present a detailed analysis process. To establish the groundwork for our proof, we first introduce an auxiliary lemma demonstrating the nonnegativity and boundedness of the system matrix. Before proving the boundedness result, we first show that $A_{Q}$ is elementwise nonnegative, which will be used in the subsequent lemma.
Lemma 4 ([13]).
For any $Q\in{\mathbb{R}}^{|{\cal S}\times{\cal A}|}$, $A_{Q}$ is a nonnegative matrix (all entries are nonnegative).
Proof.
Recalling the definition $A_{Q}:=I+\alpha(\gamma DP\Pi_{Q}-D)$, one can easily see that for any $i,j\in\{1,2,\ldots,|{\cal S}\times{\cal A}|\}$, we have ${[{A_{Q}}]_{ij}}={[I-\alpha D+\alpha\gamma DP{\Pi_{Q}}]_{ij}}={[I-\alpha D]_{ij}}+\alpha\gamma{[DP{\Pi_{Q}}]_{ij}}\geq 0$, where $\cdot]_{ij}$} denotes the element of a matrix $\cdot$] in the $i$th row and $j$th column, and the inequality follows from the fact that both $I-\alpha D$ and $DP{\Pi_{Q}}$ are nonnegative matrices. This completes the proof. ∎
Having established that $A_{Q}$ is elementwise nonnegative, we next analyze its boundedness property, which plays a crucial role in ensuring the stability of the subsequent system dynamics.
Lemma 5 ([13]).
For any $Q\in{\mathbb{R}}^{\lvert\mathcal{S}\rvert\lvert\mathcal{A}\rvert}$, we have
@@
$\begin{split}\lVert A_{Q}\rVert_{\infty}&=\max_{i\in\{1,2,...,\lvert S\rvert\lvert A\rvert\}}{1+\alpha[D]_{ii}(\gamma-1)}\\ &=1-\alpha\min_{(s,a)\in\mathcal{S}\times\mathcal{A}}d(s,a)(1-\gamma)\\ &=\rho,\end{split}$
@@ B.1 Convergence of $Q_{k}^{\mathrm{err}_{U}}$
$Q_{k+1}^{\text{err}_{U}}=(I+\alpha\gamma DP\Pi_{Q_{k}^{\text{err}_{U}}}-\alpha D)Q_{k}^{\text{err}_{U}}+\alpha w_{k}^{A}-\alpha w_{k}^{B},\quad Q_{0}^{\text{err}_{U}}\in\mathbb{R}^{\lvert\mathcal{S}\rvert\lvert\mathcal{A}\rvert}$, (19)
@@
Proposition A.1.
Suppose $Q_{0}^{\mathrm{err}_{U}}\geq Q_{0}^{\mathrm{err}}$, where $\geq$” is used as the element-wise inequality. Then, we have
$Q_{k}^{\mathrm{err}_{U}}\geq Q_{k}^{\mathrm{err}},$
@@
$\begin{split}Q_{k+1}^{\text{err}}&=Q_{k}^{\text{err}}+\alpha\gamma DP\Pi_{Q_{k}^{B}}Q_{k}^{A}-\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{B}-\alpha DQ_{k}^{\text{err}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &\leq Q_{k}^{\text{err}}+\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{A}-\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{B}-\alpha DQ_{k}^{\text{err}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &=Q_{k}^{\text{err}}+\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{\text{err}}-\alpha DQ_{k}^{\text{err}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &=(I+\alpha\gamma DP\Pi_{Q_{k}^{A}}-\alpha D)Q_{k}^{\text{err}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &\leq(I+\alpha\gamma DP\Pi_{Q_{k}^{\text{err}}}-\alpha D)Q_{k}^{\text{err}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &\leq(I+\alpha\gamma DP\Pi_{Q_{k}^{\text{err}}}-\alpha D)Q_{k}^{\text{err}_{U}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &\leq(I+\alpha\gamma DP\Pi_{Q_{k}^{\text{err}_{U}}}-\alpha D)Q_{k}^{\text{err}_{U}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &=Q_{k+1}^{\text{err}_{U}},\end{split}$
@@
To prove the convergence of $Q_{k}^{\mathrm{err}_{U}}$, we consider another comparison system $Q_{k}^{\mathrm{err}_{UL}}$ which is a lower comparison system of $Q_{k}^{\mathrm{err}_{U}}$. In the following subsection, a convergence analysis of $Q_{k}^{\mathrm{err}_{UL}}$ is presented.
B.2 Convergence of $Q_{k}^{\mathrm{err}_{UL}}$
Let us write the $Q_{k}^{\mathrm{err}_{UL}}$, which has the following form:
$Q_{k+1}^{\mathrm{err}_{UL}}=(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)Q_{k}^{\mathrm{err}_{UL}}+\alpha w_{k}^{A}-\alpha w_{k}^{B},$ (20)
where the stochastic noises $w_{k}^{A}$ and $w_{k}^{B}$ are identical to those in the original system (5.1). Note that the system is the lower comparison system of the upper comparison system corresponding to $Q_{k}^{\mathrm{err}}$. In the following proposition, we prove that $Q_{k}^{\text{err}_{UL}}$ is a lower comparison system of $Q_{k}^{\text{err}_{U}}$.
Proposition A.2.